OpenStudy (anonymous):

solutions to (x-2)^2=16? help?

OpenStudy (anonymous):

so far I have gotten x^2-4=16

OpenStudy (anonymous):

\[x=\pm4+2\]

OpenStudy (anonymous):

@gblover11...ur equation has some error...the equation is x^2-4x+4=16, check urs...

OpenStudy (anonymous):

(x-6)(x+2)=0 therefore x=6,x=-2

OpenStudy (anonymous):

you mean what I have worked out is wrong?

OpenStudy (anonymous):

yes...post ur work,we'll tell u wer u went wrong :)

OpenStudy (anonymous):

I did (x-2)^2 and got x^2-4=16